Black hair (B) is dominant to red hair (b). If there are 175 black haired people in a population of 250: (Answer in percent or decimal).


A. What is the predicted frequency of heterozygotes? 

B. What is the predicted frequency of homozygous dominant? 

C. What is the predicted frequency of homozygous recessive?





1
Expert's answer
2021-02-08T10:21:24-0500

The frequency of red-haired people:

f (bb) = (250 - 175) / 250 = 0.3

From here the frequency of red and black alleles are:

f (b) = 0.31/2 = 0.55

f (B) = 1 - f(b) = 1 - 0.55 = 0.45

A. The frequency of heterozygotes:

f (Bb) = 2 × f (B) × f (b) = 2 × 0.55 × 0.45 = 0.5

B. The predicted frequency of homozygous dominant (BB):

f (BB) = f(B)2 = 0.452 = 0.2

C. The predicted frequency of homozygous recessive (bb):

f (bb) = f(b)2 = 0.552 = 0.3